# Diffusion Model with Perceptual Loss

### v6

@@ 6.1 Layer $l$: Only Midblock Layer is Better
We compare the effect of computing loss on features from different layers $l$. We try using all the features from the encoder and decoder layers by summing up the losses but find that only using the features from the midblock layer yields better results, as shown in Tab. 3.
@@ 6.3 Feature Distance Function: Not Influential
6.4 Other Possible Formulations
We experiment with an alternative formulation, which combines predicted $\hat{x}_{t^{\prime}},\hat{\epsilon}_{t^{\prime}}$ separately with ground-truth $x_{t^{\prime}},\epsilon_{t^{\prime}}$. This formulation allows gradient feedback at $t^{\prime}=t$:
$x_{t^{\prime}}$ $=\mathbf{forward}(x_{0},\epsilon,t^{\prime}),$ (12)
$\hat{x}_{t^{\prime}}^{x}$ $=\mathbf{forward}(\hat{x}_{0},\epsilon,t^{\prime}),$ (13)
$\hat{x}_{t^{\prime}}^{\epsilon}$ $=\mathbf{forward}(x_{0},\hat{\epsilon},t^{\prime}),$ (14)
$\mathcal{L}_{sp2}=$ $\|f^{l}_{*}(\hat{x}_{t^{\prime}}^{x},t^{\prime},c)-f^{l}_{*}(x_{t^{\prime}},t^{\prime},c)\|_{2}^{2}$ (15)
$+$ $\|f^{l}_{*}(\hat{x}_{t^{\prime}}^{\epsilon},t^{\prime},c)-f^{l}_{*}(x_{t^{\prime}},t^{\prime},c)\|_{2}^{2}.$
Table 6 shows that the alternative formulation yields worse performance.
Formulation FID $\downarrow$ IS $\uparrow$
$\mathcal{L}_{sp}$ 25.89 27.76
$\mathcal{L}_{sp2}$ 29.83 24.54
Table 6: Comparing different formulations. We find that merged formulation yields the best results.
@@ 6.6 Combine with Classifier-Free Guidance
We experiment with applying classifier-free guidance on the model trained with our self-perceptual objective. Table 8 shows that classifier-free guidance indeed can improve sample quality further on the self-perceptual model but it does not surpass classifier-free guidance applied on the MSE model.
@@ 6.7 Artifacts Caused by the Perceptual Network
In Fig. 6, we visualize the model prediction by converting to $\hat{x}_{0}$ at every inference step. We notice that it has grid-like pattern artifacts, likely resulting from the perceptual network using convolution with kernel size 3 and stride 2 for downsampling [32].
@@ 6.8 Limitations and Future Works
As stated in Sec. 6.7, the convolution kernel in the existing network architecture can cause undesirable artifacts. Furthermore, prior work has shown that the perceptual network can be tuned to even better align with human preferences [54]. Using recaption networks to improve the training dataset’s text alignment may eliminate CFG’s unfair advantage on text alignment. We believe these improvements can lead to further improvements but we leave them to future works.
@@ 7 Conclusion
Our paper elucidates that the loss objective’s important role in shaping the learned distribution of diffusion models. We show that the squared distance causes the diffusion models to generate poor samples without guidance and demonstrate the effectiveness of perceptual loss in improving sample quality. We hope our work paves the way for more future explorations on diffusion training objectives.

### v7

@@ 6.1 Layer $l$: Only Midblock Layer is Better
We compare the effect of computing loss on features from different layers $l$. We find that only using the features from the midblock layer yields better results, as shown in Tab. 3.
@@ 6.5 Combine with Classifier-Free Guidance
We experiment with applying classifier-free guidance on the model trained with our self-perceptual objective. Table 7 shows that classifier-free guidance indeed can improve sample quality further on the self-perceptual model but it does not surpass classifier-free guidance applied on the MSE model. We find artifacts as described in Sec. 6.6.
@@ 6.6 Artifacts Caused by the Perceptual Network
In Fig. 6, we visualize the model prediction by converting to $\hat{x}_{0}$ at every inference step. We see grid-like pattern artifacts resulting from the perceptual network using convolution with kernel size 3 and stride 2 for downsampling [33]. This can be an area for minor future improvement.
@@ 6.7 Limitations and Future Works
In this work, we compute distance on the features of a frozen network. Although this exhibits better perceptual alignment than MSE, it is still not ideal. Our proposed SP objective is only meant to validate the effect of the loss objective and is not proposed as a final solution. We believe the loss objective should ultimately be learned. This draws a connection to adversarial training where the discriminator is a learnable perceptual network in the loop. We leave the exploration to future work.
@@ 7 Conclusion
Our paper elucidates that the loss objective has an important role in shaping the learned distribution of diffusion models. We show that the MSE loss causes the diffusion models to generate poor samples without guidance and demonstrate the effectiveness of perceptual loss in improving sample quality. We hope our work paves the way for more future explorations on diffusion training objectives.